# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ 8 Discussion
In this paper, we have established sufficient conditions of directional convergence of gradient descent in leaky ReLU two-layer neural networks and, as an application, also derived sufficient conditions of benign overfitting. Here we discuss directions for future research other than the obvious direction of extending to deeper neural networks:
Our assumptions of Theorem 6.1 and 6.2 are stronger than what was imposed in prior works on binary linear classification (Wang and Thrampoulidis (2022); Cao et al. (2021); Hashimoto et al. (2025)). This is due to the asymmetric nature of leaky ReLU activation, which leads to complications that are not present in binary linear classification. First, gradient descent iterates are more difficult to analyze. So, we had to impose additional assumptions to obtain directional convergence. Second, due to the more complicated structure of $\tilde{X}\tilde{X}^{\top}$, we had to base our analysis of $(\tilde{X}\tilde{X}^{\top})^{-1}$ on cruder techniques than prior work on binary linear classification. Given prior work on gradient flow (Lyu and Li (2020); Ji and Telgarsky (2020)), we conjecture that directional convergence can be shown only assuming linear separability of the training data, which is the only assumption necessary for binary linear classification (Soudry et al. (2018)).
Our analysis revealed that the equivalence of $\boldsymbol{\hat{w}}$ and the minimum norm least square estimator fails in the strong signal regime if $q_{\gamma}<\gamma$. We conjecture that $\boldsymbol{\hat{w}}$ would be equivalent to the minimum norm least square only using positive-labeled samples (or negative-labeled samples) in this case.
@@
In this study, we did not introduce noise to the label $y$ as was done in some prior works on binary classification (Chatterji and Long (2021); Wang and Thrampoulidis (2022); Frei et al. (2022); Xu and Gu (2023); Hashimoto et al. (2025)). Introducing noise to the label would likely not change the result for the weak signal regime since in this regime the signal $\boldsymbol{\mu}$ is negligible. However, we conjecture that the convergent direction would be very different in the strong signal regime with or without noise, as was revealed in Hashimoto et al. (2025) for binary linear classification, and hence require more involved argument.
Appendix A Proof of Directional Convergence
A.1 Proof of Theorem 4.1
Lemma A.1.
@@ Appendix A Proof of Directional Convergence
Lemma A.1 implies that the gradient descent update is same within $j\in J_{+}$ and $j\in J_{-}$, respectively.
Lemma A.2.
@@
$0>\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle=\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\geq\frac{\gamma|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle=f(\boldsymbol{x};\hat{W}).$
@@
With this, the gradient descent updates in Lemma A.1 can be rewritten as
@@
Now we turn to the characterization of the convergent direction. First, we show that not only $\|\boldsymbol{w}^{(t)}\|$ but also $\|\boldsymbol{w}_{j}^{(t)}\|$ diverges to infinity for all $j\in J$. To see this, note that by Lemma A.1, we have for $t\geq T$
@@
Again by Lemma A.1, we have for any $j$,
@@
Finally, the claim about the linear decision boundary follows from Lemma A.2
Before proceeding to the subsequent subsections where we prove Theorem 4.4 and 4.9, we introduce some ancillary results:
@@
A.2 Proof of Theorem 4.4
Lemma A.4.
@@
×$\qquad\qquad\qquad\qquad\times\exp\{-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\}-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\Big{]}$
@@
Proof of Theorem 4.4.
By Lemma A.4, all the neurons stay activated after $t\geq 1$.
@@
By Lemma A.3 and $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ we have
$\sigma_{max}^{2}(\tilde{X})$ $\leq(1+\theta_{2})(\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z}))+2(n-1)\|\boldsymbol{\mu}\|^{2}$ (25)
@@
Therefore, the directional convergence follows by Theorem 4.1. ∎
A.3 Proof of Theorem 4.9
@@
Lemma A.5.
Suppose event $E$, Assumption 4.5 and 4.6 with $\varepsilon_{1}+\varepsilon_{2}<1$. Then, we have for any $i,j$,
@@
Lemma A.6.
Suppose event $E$, Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$. Then we have
@@
where the second inequality is due to Assumption 4.5 and 4.7, the third is due to Assumption 4.6, the fourth is due to $\varepsilon_{1}+\varepsilon_{2}<1$, and the last inequality is due to Assumption 4.8.
Therefore, with Lemma A.5, we have $0<y_{i}f(\boldsymbol{x}_{i};W^{(1)})\leq 1$, and hence
@@
We now show $\alpha\sigma_{max}^{2}(\tilde{X})<2$. By Lemma A.3, we have
$\alpha\sigma_{max}^{2}(\tilde{X})$ $\leq\alpha\left\{(1+\theta_{2})\{n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}+\varepsilon_{1}(1+\theta_{2})R_{max}^{2}(\boldsymbol{z})\right\}$
@@
Lemma A.7.
Suppose event $E$, Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}<\frac{1}{2},i=1,3$, and $\theta_{2}\leq\frac{1}{2}$. Then we have for any $i,j$
@@
We prove both claims simultaneously by induction. By Lemma A.5 and A.6, both claims hold at step $t=1$. Suppose both claims hold for $t\geq 1$.
@@
×$\quad\times\Big{\{}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})-C_{w}\varepsilon_{1}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\Big{\}}$
@@
×$\quad\times\Big{[}(1+\theta_{2})\big{(}1+\frac{\varepsilon_{3}}{n}\big{)}R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{3}R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{1}(1+\theta_{2})R_{max}^{2}(\boldsymbol{z})\Big{]}$
@@
where the first inequality is due to Assumption 4.5 and 4.7, the last is due to the assumptions on $\varepsilon_{i}$.
@@
Proof of Theorem 4.9.
The conclusion now follows immediately from Lemma A.6, A.7, and Theorem 4.1. ∎
Appendix B Proof of Theorem 5.1
@@
To obtain approximation of $(\tilde{X}\tilde{X}^{\top})^{-1}$ from Lemma B.3, we need the following ancillary lemma:
Lemma B.4.
@@
Note that the additional assumption implies $\tilde{\varepsilon}q_{\gamma}^{-1}\leq\tfrac{1}{2}$. Then, the desired conclusion follows from Lemma B.3 and B.4 by taking $U=B^{-1}\tilde{X}\tilde{X}^{\top}B^{-1},V=A,s=\tilde{\varepsilon}q_{\gamma}^{-1}$, and $L=R$. ∎
The next lemma establishes sufficient conditions of $\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}y$. Before proceeding to the lemma, we note that the following inequality holds under $\tilde{E}_{3}$:
Lemma B.6.
  1. (i)
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{\big{\{}(1+\tilde{\varepsilon}_{3})\lambda+2\big{\}}^{2}}$ for some $\lambda<\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{(1-\tilde{\varepsilon}_{3})q_{\gamma}^{2}}{4}\tilde{C}(\gamma,q_{+},\tilde{\varepsilon}_{3},\lambda)\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$ for some $\lambda>0$, where
    $\tilde{C}(\gamma,q_{+},\lambda):=\frac{q_{\gamma}-\gamma}{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}}.$
Then, we have
$\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}y.$
@@
By Lemma B.1, it suffices to show $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>0$ for all $i$.
By Lemma B.5, we have
where the last inequality follows from the definition of $q_{\gamma}$ and $n_{+}+n_{-}=n$.
@@
Let $y=\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$, where $(\boldsymbol{1}_{+})_{i}=1$ if $i\in I_{+}$ and $(\boldsymbol{1}_{+})_{i}=0$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=1$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=0$ if $i\in I_{+}$.
@@
Thus, we have
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{+}^{-1}y_{i}\boldsymbol{e}_{i}$ (37)
$=$ $\begin{cases}R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right),&\text{if $i\in I_{+}$},\\ \gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, by the definition of $\tilde{E}_{3}$, (34) and (37), we have
$\geq$ $\begin{cases}R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{+}$},\\ \gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{-}$}.\end{cases}$
@@
Thus, we have
$(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{-}^{-1}y_{i}\boldsymbol{e}_{i}$ (40)
$=$ $\begin{cases}\gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right),&\text{if $i\in I_{+}$}\\ R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, the definition of $\tilde{E}_{3}$, (34) and (40), we have
$\geq$ $\begin{cases}\gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{+}$},\\ R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, for both the cases above, it suffices to have
which follows from the condition (i).
@@
×$\quad\times\Big{[}\sqrt{q_{+}+\gamma^{2}q_{-}}\big{\{}(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big{\}}\boldsymbol{1}_{+}$
@@
Therefore, by (34) and (45) , we have for $i\in I_{+}$
@@
×$\qquad\times\Big{\{}\sqrt{q_{+}+\gamma^{2}q_{-}}\big{\{}(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big{\}}$

### v2

@@ 8 Discussion
In this paper, we have established sufficient conditions of benign overfitting of gradient descent in leaky ReLU two-layer neural networks. Here we discuss directions for future research other than the obvious direction of extending to deep neural networks:
Not only showing directional convergence in more generality but also obtaining detailed characterization of the convergent direction remains an important open question. Even for two-layer neural networks we studied, our setting is rather restricted considering a more general result obtained by Ji and Telgarsky (2020) for gradient flow.
@@
In this study, we did not introduce noise to the label $y$ as was done in some prior work on binary classification (Chatterji and Long (2021); Wang and Thrampoulidis (2022); Frei et al. (2022); Xu and Gu (2023); Hashimoto et al. (2025)). We conjecture that introducing noise will not change the result for the weak signal regime since in this regime the signal $\boldsymbol{\mu}$ is negligible. However, the situation is likely to change substantially in the strong signal regime with noisy data as was observed in Hashimoto et al. (2025) for binary linear classification.
Appendix A Comparison with the existing results
In this section, we provide more detailed comparison of our main results with the existing results (Frei et al. [2022], Xu and Gu [2023]). Since Frei et al. [2022] only considers smoothed approximation of leaky ReLU, our comparison is mostly with Xu and Gu [2023].
Since only isotropic sub-Gaussian mixture was studied in the existing literature, we present here sufficient conditions of benign overfitting derived from Theorem 6.2 in the case of $\Sigma=I_{P}$ for the purpose of direct comparison:
Corollary A.1.
Consider model (sG) with $\Sigma=I_{p}$. Then, for sufficiently large $n$, benign overfitting occurs with the convergent direction $\hat{W}$ with high probability if either one of the following holds:
  1. (i)
  2. (ii)
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and $n^{\frac{5}{4}}\|\boldsymbol{\mu}\|^{\frac{3}{2}}\lesssim p\lesssim\min\{n\|\boldsymbol{\mu}\|^{2},\tfrac{\|\boldsymbol{\mu}\|^{4}}{n}\}$,
  3. (iii)
We first note that conditions (i) and (ii) are newly identified regimes on which benign overfitting occurs, and only condition (iii) is directly comparable to the results studied in the prior work (Frei et al. [2022], Xu and Gu [2023]). As shown in Theorem 6.3, we have further extended the result in the case of (PM).
Data Regime: Our work considers both weak and strong signal regimes and identified the surprising phase transition in classification error occurring at $n\|\boldsymbol{\mu}\|^{2}\approx{\rm tr}(\Sigma)$. Frei et al. [2022], Xu and Gu [2023] only studied nearly orthogonal data regime, which corresponds to the weak signal regime in this paper, and thus did not identify the phase transition. Even in the weak signal regime, condition (i) in Corollary A.1 was not studied.
Distributional setting: Our results significantly relax distributional conditions in prior work by extending to anisotropic polynomially tailed mixture (PM). Both Frei et al. [2022], Xu and Gu [2023] considered mixtures even stronger than (sG). Specifically, Frei et al. [2022] assumed $\boldsymbol{z}$ is not only sub-Gaussian but also strongly log-concave. Xu and Gu [2023] assumed $\boldsymbol{z}$ has a bounded log-Sobolev constant, which is also stronger than just assuming $\boldsymbol{z}$ is sub-Gaussian. Moreover, both work only considered isotropic mixture.
Comparing condition (iii) with that of Xu and Gu [2023], we have a better upper bound on $p$ (Xu and Gu [2023] requires $p\log(mp)\lesssim n\|\boldsymbol{\mu}\|^{4}$). We allow the width $m$ to be arbitrary while Xu and Gu [2023] requires $m\gtrsim\log n$. We note that Xu and Gu [2023]’s result holds for non-smooth leaky ReLU like activation, which is not limited to (leaky) ReLU. We also note that our lower bound on $p$ is slightly stronger than theirs which requires $p\gtrsim\max\{n\|\boldsymbol{\mu}\|^{2},n^{2}\log n\}$.
Appendix B Proof of Directional Convergence
B.1 Proof of Lemma 7.1
Lemma B.1.
@@ Appendix B Proof of Directional Convergence
Lemma B.1 implies that the gradient descent update is same within $j\in J_{+}$ and $j\in J_{-}$, respectively.
Lemma B.2.
@@
[…]
@@
With this, the gradient descent updates in Lemma B.1 can be rewritten as
@@
Now we turn to the characterization of the convergent direction. First, we show that not only $\|\boldsymbol{w}^{(t)}\|$ but also $\|\boldsymbol{w}_{j}^{(t)}\|$ diverges to infinity for all $j\in J$. To see this, note that by Lemma B.1, we have for $t\geq T$
@@
Again by Lemma B.1, we have for any $j$,
@@
Finally, the claim about the linear decision boundary follows from Lemma B.2
Before proceeding to the subsequent subsections where we prove Theorem 4.8, we introduce some ancillary results:
@@
B.2 Proof of Theorem 4.8 under condition (i)
Lemma B.4.
@@
$\qquad\qquad\qquad\qquad\times\exp\{-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\}-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\Big]$
@@
Proof of Theorem 4.8 under condition (i).
By Lemma B.4, all the neurons stay activated after $t\geq 1$.
@@
By Lemma B.3 and $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ we have
[…]
@@
Therefore, the directional convergence follows by Lemma 7.1. ∎
B.3 Proof of Theorem 4.8 under condition (ii)
@@
Lemma B.5.
Suppose event $E$, Assumption 4.4 and 4.5 with $\varepsilon_{1}+\varepsilon_{2}<1$. Then, we have for any $i,j$,
@@
Lemma B.6.
Suppose event $E$, Assumption 4.4, 4.5, 4.6, and 4.7 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$. Then we have
@@
where the second inequality is due to Assumption 4.4 and 4.6, the third is due to Assumption 4.5, the fourth is due to $\varepsilon_{1}+\varepsilon_{2}<1$, and the last inequality is due to Assumption 4.7.
Therefore, with Lemma B.5, we have $0<y_{i}f(\boldsymbol{x}_{i};W^{(1)})\leq 1$, and hence
@@
We now show $\alpha\sigma_{max}^{2}(\tilde{X})<2$. By Lemma B.3, we have
[…]
@@
Lemma B.7.
Suppose event $E$, Assumption 4.4, 4.5, 4.6, and 4.7 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}<\frac{1}{2},i=1,3$, and $\theta_{2}\leq\frac{1}{2}$. Then we have for any $i,j$
@@
We prove both claims simultaneously by induction. By Lemma B.5 and B.6, both claims hold at step $t=1$. Suppose both claims hold for $t\geq 1$.
@@
$\quad\times\Big\{\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})-C_{w}\varepsilon_{1}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\Big\}$
@@
$\quad\times\Big[(1+\theta_{2})\big(1+\frac{\varepsilon_{3}}{n}\big)R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{3}R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{1}(1+\theta_{2})R_{max}^{2}(\boldsymbol{z})\Big]$
@@
where the first inequality is due to Assumption 4.4 and 4.6, the last is due to the assumptions on $\varepsilon_{i}$.
@@
Proof of Theorem 4.8 under condition (ii).
The conclusion now follows immediately from Lemma B.6, B.7, and Lemma 7.1. ∎
Appendix C Proof of Theorem 5.1
@@
To obtain approximation of $(\tilde{X}\tilde{X}^{\top})^{-1}$ from Lemma C.3, we need the following ancillary lemma:
Lemma C.4.
@@
Note that the additional assumption implies $\tilde{\varepsilon}q_{\gamma}^{-1}\leq\tfrac{1}{2}$. Then, the desired conclusion follows from Lemma C.3 and C.4 by taking $U=B^{-1}\tilde{X}\tilde{X}^{\top}B^{-1},V=A,s=\tilde{\varepsilon}q_{\gamma}^{-1}$, and $L=R$. ∎
By Lemma C.1 and C.5, the rest of our arguments is reduced to the analysis of $(B^{-1}\boldsymbol{y})^{\top}A^{-1}$. Here, we present technical results regarding this quantity which are used frequently later.
Lemma C.6.
Let $\boldsymbol{y}=\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$, where $(\boldsymbol{1}_{+})_{i}=1$ if $i\in I_{+}$ and $(\boldsymbol{1}_{+})_{i}=0$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=1$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=0$ if $i\in I_{+}$. We have the following equalities:
  1. 1.
    if $q_{+}=1,q_{-}=0$,
    $(B^{-1}\boldsymbol{y})^{\top}A^{-1}=$ $R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)\boldsymbol{1}_{+}\right$. (33)
    $\qquad\qquad\left.-\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)\boldsymbol{1}_{-}\right\}^{\top},$
  2. 2.
    if $q_{+}=0,q_{-}=1$,
    $(B^{-1}\boldsymbol{y})^{\top}A^{-1}=$ $R^{-1}\left\{\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right)\boldsymbol{1}_{+}\right$. (34)
    $\qquad\qquad\left.-\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right)\boldsymbol{1}_{-}\right\}^{\top},$
  3. 3.
    if $q_{+}q_{-}\neq 0$,
    $(B^{-1}\boldsymbol{y})^{\top}A^{-1}$ (35)
    $=$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
    $\quad\times\Big[\sqrt{q_{+}+\gamma^{2}q_{-}}\big\{(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big\}\boldsymbol{1}_{+}$
    $\qquad-\sqrt{q_{-}+\gamma^{2}q_{+}}\big\{(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big\}\boldsymbol{1}_{-}\Big]^{\top},$
    where
    $d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ (36)
    $\qquad=(1-\gamma^{2})^{2}q_{+}q_{-}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}+(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n\|\boldsymbol{\mu}\|^{2}+R\}R.$
@@
$\quad\times\Big[\sqrt{q_{+}+\gamma^{2}q_{-}}\big\{(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big\}\boldsymbol{1}_{+}$
@@
Before proceeding to the next lemma, we note that the following inequality holds under $\tilde{E}_{3}$:
The next lemma establishes useful bounds on $d$.
Lemma C.7.
We have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ $\leq\left\{\gamma^{2}+\frac{5(1-\gamma^{2})^{2}}{4}q_{+}q_{-}\right\}\left(n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}$ (43)
$\leq\frac{5q_{\gamma}}{4}\left(n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}.$
If additionally $\tilde{E}_{3}(\tilde{\varepsilon}_{3})$ holds for some $\tilde{\varepsilon}_{3}\in[0,1]$, then we also have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\geq\frac{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}{8}\left(n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}+\gamma^{2}R^{2}.$ (44)
Proof.
We first note that
$q_{\gamma}\geq(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})=(1+\gamma^{4})q_{+}q_{-}+\gamma^{2}(q_{+}^{2}+q_{-}^{2})=\gamma^{2}+(1-\gamma^{2})^{2}q_{+}q_{-}.$ (45)
With (45) and also noting that $n_{+}n_{-}\leq\frac{n^{2}}{4}$, it is easy to see (43) follows from (36).
By (36), (45), and (42), we have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$
$\geq$ $\frac{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}}{4}q_{+}q_{-}\left(n\|\boldsymbol{\mu}\|^{2}\right)^{2}+\left\{\gamma^{2}+(1-\gamma^{2})^{2}q_{+}q_{-}\right\}R^{2}$
$\geq$ $\frac{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}}{4}q_{+}q_{-}\left\{\left(n\|\boldsymbol{\mu}\|^{2}\right)^{2}+R^{2}\right\}+\gamma^{2}R^{2}$
$\geq$ $\frac{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}}{8}q_{+}q_{-}\left(n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}+\gamma^{2}R^{2},$
where the last inequality is due to $a^{2}+b^{2}\geq\frac{(a+b)^{2}}{2}$.
Now we are ready to establish sufficient conditions of $\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$.
Lemma C.8.
  1. (i)
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{5(1+\lambda)^{2}}$ for some $\lambda\leq\frac{q_{\gamma}}{2\gamma(1-\gamma)}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\tilde{C}(\gamma,q_{+},\tilde{\varepsilon}_{3},\lambda)\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$ for some $\lambda>0$, where
    $\tilde{C}(\gamma,q_{+},\tilde{\varepsilon}_{3},\lambda):=\frac{(1-\tilde{\varepsilon}_{3})q_{\gamma}^{2}(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})^{2}}.$
Then, we have
$\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}y.$
Proof.
By Lemma C.1, it suffices to show $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>0$ for all $i$.
By Lemma C.5, we have
where the last inequality follows from the definition of $q_{\gamma}$ and $n_{+}+n_{-}=n$. The rest of the argument is split in cases.
Case $q_{+}=1,q_{-}=0$: By (33), we have
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{+}^{-1}y_{i}\boldsymbol{e}_{i}$ (47)
[…]
Therefore, by the definition of $\tilde{E}_{3}$, (46) and (47), we have
[…]
Case $q_{+}=0,q_{-}=1$: By (34), we have
$(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{-}^{-1}y_{i}\boldsymbol{e}_{i}$ (49)
[…]
Therefore, by the definition of $\tilde{E}_{3}$, (46) and (49), we have
[…]
Therefore, for both the cases above, it suffices to have
which follows from the condition (i).
Case $q_{+}q_{-}\neq 0$: By (41) and (46), we have for $i\in I_{+}$
@@
$\qquad\times\Big\{\sqrt{q_{+}+\gamma^{2}q_{-}}\big\{(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big\}$